# Rethinking Visual Token Reduction in LVLMs under Cross-modal Misalignment

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) encode visual inputs as dense sequences of patch-level tokens to capture fine-grained semantics. These visual tokens often outnumber their textual counterparts by a large margin, leading to substantial computational overhead and limiting the scalability of LVLMs in practice. Previous efforts have explored visual token reduction either prior to or within the large language models (LLM). However, most in-LLM reduction approaches rely on text-conditioned interactions, implicitly assuming that textual tokens can reliably capture the importance of visual tokens. In this work, we revisit this assumption and reveal causal, semantic, and spatial forms of cross-modal misalignment. These misalignments undermine the effectiveness of text-guided visual token reduction. To address this, we introduce VisionDrop, a training-free, visual-only pruning framework that selects informative visual tokens based on intra-modal (visual-to-visual) attention, without relying on textual signals. To further suppress redundancy throughout the model hierarchy, we treat the visual encoder and the LLM as a unified system and design a progressive pruning pipeline. Our method performs dominant token selection and lightweight contextual merging at multiple stages, enabling fine-grained visual information to be retained even under aggressive token budgets. Extensive experiments across diverse benchmarks show that VisionDrop achieves consistent improvements over existing methods, despite requiring no additional training or complex modifications. Its simple yet effective design enables efficient inference while preserving strong performance across tasks.
@@ Introduction
Figure 1: Comparison of visual token pruning strategies based on attention maps from the LLM decoding layers. Here, $x_{V1}$ and $x_{V2}$ denote visual tokens, $x_{T1}$ and $x_{T2}$ represent text instructions, and $y_{1}$, $y_{2}$ are autoregressively generated output tokens. (a) Our method identifies important visual tokens (e.g., the trees highlighted in orange) by leveraging image-to-image attention (blue box), which reflects intra-modal relevance and avoids interference from misaligned cross-modal signals. (b) In contrast, previous methods rely on image-to-text attention (red box) to assess visual token importance, which can be overly sensitive to cross-modal noise, resulting in the preservation of semantically redundant visual regions (e.g., the grass in green).
To address this inefficiency, recent studies have explored reducing the number of visual tokens either before or within the layers of the large language models (LLM) (Yang et al. 2025; Zhang et al. 2025c; Chen et al. 2024b; Xing et al. 2025). Pruning visual tokens before entering the LLM enables effective compression of redundant features at the representation level, but may risk discarding subtle or task-relevant details. In contrast, performing token reduction within the LLM allows uninformative tokens to be identified after visual information has been integrated into the language context, though it risks discarding entangled or contextually important features. Additionally, a more fundamental limitation of many in-LLM pruning strategies lies in their reliance on text-conditioned scoring mechanisms, which estimate the relevance between visual tokens and textual inputs. These approaches implicitly assume strong and persistent alignment between modalities throughout the LLM layers.
@@
Based on the analysis, we propose VisionDrop, a training-free framework for visual token reduction. Unlike prior approaches that rely on text-guided relevance, VisionDrop estimates token importance solely from visual self-attention, avoiding dependence on potentially misaligned textual cues. Our method performs stage-wise pruning across the full LVLM architecture, progressively reducing visual tokens in both the visual encoder and LLM layers. First, the dominant token selection identifies highly referenced visual tokens by the aforementioned visual-only importance score, ensuring that key semantic content is retained. Second, the lightweight contextual merging aggregates remaining tokens into contextual tokens by similarity, preserving complementary information. Together, these components are applied at multiple stages in the model, allowing VisionDrop to maintain expressive visual representations even under tight token budgets.
@@
Extensive evaluations on nine standard benchmarks demonstrate that VisionDrop achieves superior performance compared to state-of-the-art token reduction methods.
@@
The evolution of large vision-language models (LVLMs) has been driven by the impressive generalization and reasoning abilities of large language models (LLMs) (Touvron et al. 2023; Jiang et al. 2023; Bai et al. 2023; Qwen et al. 2025; Cai et al. 2024), extending their success to the multimodal domain. Recent LVLMs integrate visual and textual modalities by projecting visual inputs into token sequences compatible with language models (Liu et al. 2023; Li et al. 2024; Wang et al. 2024; Bai et al. 2025; Chen et al. 2025b; Zhu et al. 2025), enabling them to leverage the full capacity of LLMs. However, the information density in visual data is often much lower than in textual data, resulting in an excessive number of visual tokens. For instance, LLaVA-1.5 (Liu et al. 2024a) encodes a $336\times 336$ image into 576 tokens, while LLaVA-NeXT (Liu et al. 2024b) expands this to 2,880 tokens for high-resolution input. This issue becomes even more critical in video understanding tasks. For example, LongVA (Zhang et al. 2024) transforms 2,000 frames into more than 200K visual tokens, while LongVILA (Chen et al. 2025a) supports inputs exceeding 6,000 frames and produces over a million visual tokens. The inflated visual token sequences place heavy demands on computation and memory, motivating the need for effective token reduction strategies to improve scalability and efficiency across tasks.
@@
Figure 3: Visualization of text-guided visual token retention after shallow-layer pruning (specifically, the second layer) in the LLM. Pale translucent blocks indicate pruned tokens. Retained tokens consistently cluster at the bottom of the image, revealing a spatial bias caused by the causal attention in autoregressive LLMs.
@@
To mitigate the computational burden in LVLMs, recent studies have explored reducing the number of visual tokens. Existing approaches to visual token reduction are primarily applied either in the visual encoder or within the LLM. They typically use either text-guided or text-agnostic scoring strategies to determine token importance. Visual encoder pruning approaches perform token reduction before tokens are fed into the LLM. Some approaches rely solely on visual information: FlowCut (Tong et al. 2025) leverages token-level information flow across layers to identify and prune redundant visual tokens in a progressive and structure-aware manner. VisionZip (Yang et al. 2025) selects dominant tokens based on attention scores and merges the rest using semantic similarity. VScan (Zhang et al. 2025a) introduces a two-stage reduction pipeline by applying global-local scanning in the visual encoder pruning. Other works utilize textual guidance at this stage. CDPruner (Zhang et al. 2025b) maximizes instruction-conditioned diversity of retained tokens. SparseVLM (Zhang et al. 2025c) computes token importance via cross-modal attention with text guidance. Although visual encoder pruning benefits from more structured token semantics, it may risk losing fine-grained cues critical for multimodal reasoning. LLM decoding pruning is an alternative direction that performs token reduction within the LLM, where visual and textual tokens are jointly processed. Most of these approaches adopt text-guided token selection. FastV (Chen et al. 2024b) ranks tokens based on their attention from text during generation, and VScan also extends its pruning into LLM layers based on cross-modal attentions. PyramidDrop (Xing et al. 2025) selects visual tokens that are most relevant to the final instruction token, performing progressive pruning across LLM layers. However, these approaches rely on the assumption that visual and textual tokens remain semantically aligned within the LLM, a property that is not guaranteed due to modality entanglement introduced by joint self-attention.
@@
To alleviate the computational overhead in LVLMs, recent studies have explored visual token reduction, either in the visual encoder (Yang et al. 2025; Zhang et al. 2025c) or during the LLM decoding phase (Chen et al. 2024b; Xing et al. 2025). In the latter case, mainstream approaches typically adopt text-guided strategies, using text relevance scores to determine which visual tokens to retain. However, this raises a critical question: Are visual and textual representations truly well-aligned within the LLM layers? Although visual inputs are projected into the textual embedding space before entering the LLM, using text-guided strategies for visual token reduction introduces three forms of cross-modal misalignment, due to the modality-specific biases in how visual and textual information are processed.
@@
In text-guided pruning, the last text instruction token is commonly used to assess the importance of visual tokens. However, due to the causal nature of autoregressive LLMs, this token tends to focus disproportionately on later-positioned visual tokens in the input sequence (Zhang et al. 2025a). Unlike the [CLS] token or average-pooled visual features in bidirectional encoders, the instruction token does not serve as an explicit global aggregator. Figure 3 demonstrates this issue, where shallow-layer pruning results show that the retained visual tokens consistently cluster near the end of the input sequence, regardless of their semantic relevance.
@@
Visual token reduction in LVLMs has been explored at two primary stages: either within the visual encoder (Yang et al. 2025; Zhang et al. 2025c) or during the LLM decoding phase (Chen et al. 2024b; Xing et al. 2025). We observe that pruning at the encoder stage tends to be more stable, as visual tokens remain semantically coherent and well-structured. However, this may lead to the loss of fine-grained visual details that are crucial for downstream reasoning. On the other hand, pruning in the LLM benefits from earlier visual-textual integration, where key visual cues may already be embedded in textual tokens, allowing more informed reduction. Yet as tokens propagate through the LLM layers, visual information becomes increasingly entangled and diffused, making it difficult to reliably estimate token importance.
To address the cross-modal misalignment within LLMs and to leverage the complementary strengths of different pruning stages, we propose a training-free, visual-only pruning framework that performs progressive visual token reduction at both the encoder and LLM stages, balancing stability and expressiveness across the model hierarchy.
@@
We partition the LVLM architecture into a sequence of stages $\mathcal{S}={s_{0},s_{1},\dots,s_{N}}$, including both the visual encoder and layers in the LLM decoder. At the end of each stage $s_{n}$, where $n=1,2,\dots,N$, we apply a stage-specific pruning ratio $\lambda_{n}$ to the current set of visual tokens $\mathbf{x}_{V}^{(s_{n})}$, where $|\mathbf{x}_{V}^{(s_{n})}|=\lambda_{n}\cdot|\mathbf{x}_{V}|$. Tokens are ranked based on visual-only attention scores, and the top-ranked subset is propagated to the next stage $s_{n+1}$. The selection is given by:
@@
To identify the most informative visual tokens, we compute attention-based importance scores by reusing the model’s self-attention maps. Specifically, we measure how frequently each token is attended to by visual query tokens, averaged over all query positions. For clarity, we present the single-head attention formulation and omit the pruning stage notation.
@@
Table 1: Performance comparison of vision-language models under different token retention rates.
@@
ToMe (ICLR23) 54.3 60.5 - 1563 72.4 65.2 68.0 52.1 - 88.49%
FastV (ECCV24) 52.7 61.2 57.0 1612 64.8 67.3 67.1 52.5 50.8 90.44%
PDrop* (CVPR25) 57.20 63.51 56.22 1783 82.31 69.56 75.59 56.1 48.9 96.61%
VisionZip* (CVPR25) 59.25 64.46 57.29 1767 86.39 68.86 76.78 57.26 51.55 98.64%
SparseVLM (ICML25) 59.5 64.1 - 1787 85.3 68.7 - 57.8 - 98.11%
VisionDrop (Ours) 60.05 65.02 58.86 1788 86.83 68.52 77.21 57.65 50.15 99.14%
@@
ToMe (ICLR23) 52.4 53.3 - 1343 62.8 59.6 63.0 49.1 - 80.38%
FastV (ECCV24) 49.6 56.1 56.4 1490 59.6 60.2 61.8 50.6 51.3 85.36%
PDrop* (CVPR25) 55.65 62.39 54.71 1662 82.01 69.66 74.21 55.14 48.49 94.64%
VisionZip* (CVPR25) 57.62 63.40 56.67 1768 84.69 68.82 75.6 56.81 52.02 97.68%
SparseVLM (ICML25) 58.4 64.5 - 1746 85.0 68.6 - 56.7 - 97.15%
VisionDrop (Ours) 58.65 65.02 56.84 1770 85.88 68.72 76.09 56.89 51.2 98.23%
@@
ToMe (ICLR23) 48.6 43.7 - 1138 52.5 50.0 57.1 45.3 - 70.11%
FastV (ECCV24) 46.1 48.0 52.7 1256 48.0 51.1 55.0 47.8 50.8 77.24%
PDrop* (CVPR25) 49.09 51.91 41.87 1427 69.75 68.27 63.7 49.56 50.66 83.92%
VisionZip* (CVPR25) 55.16 62.61 54.20 1687 80.45 69.01 72.42 55.48 52.94 95.13%
SparseVLM (ICML25) 53.8 60.1 - 1589 77.5 69.8 - 53.4 - 91.26%
VisionDrop (Ours) 55.89 62.95 55.10 1698 81.58 69.31 73.16 55.59 52.28 95.73%
@@
We apply our proposed VisionDrop to the widely adopted LLaVA-1.5-7B (Liu et al. 2024b) model, and compare our method against several state-of-the-art visual token reduction baselines, including ToMe (Bolya et al. 2023), FastV (Chen et al. 2024b), PyramidDrop (Xing et al. 2025), VisionZip (Yang et al. 2025), and SparseVLM (Zhang et al. 2025c).
@@
We evaluate our method on 9 standard image understanding benchmarks. These include general-purpose visual question answering datasets such as GQA (Hudson and Manning 2019), VQAv2 (Goyal et al. 2017), and TextVQA (Singh et al. 2019); robustness and fairness benchmarks including POPE (Li et al. 2023), MMBench (Liu et al. 2024c), MMBench-CN (Liu et al. 2024c), and MME (Fu et al. 2024); as well as specialized datasets like VizWiz (Gurari et al. 2018), which features low-quality images captured by visually impaired users, and ScienceQA (SQA) (Lu et al. 2022), which focuses on scientific reasoning over visual inputs.
@@
We adopt the standard inference settings provided in the official implementations of each LVLM. For progressive visual token pruning, we divide the model into five stages: the first stage concludes at the output of the visual encoder, while the remaining four end at decoding layers $l=8$, $16$, $24$, and the final decoding layer of the LLM. The number of visual tokens is gradually reduced across these stages. In the first stage, all original visual tokens are retained. For each target configuration—corresponding to a final token count of 192, 128, or 64—we set the retention number in the second stage to be 1.5×1.5$\times$1.5 × the final count, and prune proportionally in subsequent stages to meet the overall budget.
Results.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) encode visual inputs as dense sequences of patch-level tokens to capture fine-grained semantics. These visual tokens often outnumber their textual counterparts by a large margin, leading to substantial computational overhead and limiting the scalability of LVLMs in practice. Previous efforts have explored visual token reduction either prior to or within the large language models (LLMs). However, most in-LLM reduction approaches rely on text-conditioned interactions, implicitly assuming that textual tokens can reliably capture the importance of visual tokens. In this work, we revisit this assumption and reveal causal, semantic, and spatial forms of cross-modal misalignment. These misalignments undermine the effectiveness of text-guided visual token reduction. To address this, we introduce VisionDrop, a training-free, visual-only pruning framework that selects informative visual tokens based on intra-modal (visual-to-visual) attention, without relying on textual signals. To further suppress redundancy throughout the model hierarchy, we treat the visual encoder and the LLM as a unified system and design a progressive pruning pipeline. Our method performs dominant token selection and lightweight contextual merging at multiple stages, enabling fine-grained visual information to be retained even under aggressive token budgets. Extensive experiments across diverse benchmarks show that VisionDrop achieves consistent improvements over existing approaches, despite requiring no additional training or complex modifications. Notably, when integrated with LLaVA-NeXT-7B, VisionDrop achieves a 2.7×2.7$\times$2.7 × reduction in inference latency and 6×6$\times$6 × in FLOPs, while retaining 95.71% of the original performance.
@@ Introduction
Figure 1: Comparison of visual token pruning strategies based on attention maps from the LLM decoding layers. Here, $x_{V1}$ and $x_{V2}$ denote visual tokens, $x_{T1}$ and $x_{T2}$ represent text instructions, and $y_{1}$, $y_{2}$ are autoregressively generated output tokens. (a) Our method identifies important visual tokens (e.g., the trees highlighted in orange) by leveraging image-to-image attention (blue box), which reflects intra-modal relevance and avoids interference from misaligned cross-modal signals. (b) In contrast, previous approaches rely on image-to-text attention (red box) to assess visual token importance, which can be overly sensitive to cross-modal noise, resulting in the preservation of semantically redundant visual regions (e.g., the grass in green).
To address this inefficiency, recent studies have explored reducing the number of visual tokens either before or within the layers of the large language models (LLMs) (Yang et al. 2025; Zhang et al. 2025d; Chen et al. 2024b; Xing et al. 2025). Pruning visual tokens before entering the LLM enables effective compression of redundant features at the representation level, but may risk discarding subtle or task-relevant details. In contrast, performing token reduction within the LLM allows uninformative tokens to be identified after visual information has been integrated into the language context, though it risks discarding entangled or contextually important features. Additionally, a more fundamental limitation of many in-LLM pruning strategies lies in their reliance on text-conditioned scoring mechanisms, which estimate the relevance between visual tokens and textual inputs. They implicitly assume strong and persistent alignment between modalities throughout LLM layers.
@@
Based on the analysis, we propose VisionDrop, a training-free framework for visual token reduction. Unlike prior approaches that rely on text-guided relevance, VisionDrop estimates token importance solely from visual self-attention, avoiding dependence on potentially misaligned textual cues. Our method performs stage-wise pruning across the full LVLM architecture, progressively reducing visual tokens in both the visual encoder and LLM layers. First, the dominant token selection identifies highly referenced visual tokens by the aforementioned visual-only importance score, ensuring that key semantic content is retained. Second, the lightweight contextual merging aggregates remaining tokens into contextual tokens by similarity, preserving complementary information. Together, these components are applied at multiple stages in the model, allowing VisionDrop to retain expressive visual representations under tight token budgets.
@@
Extensive evaluations on diverse benchmarks show that VisionDrop outperforms prior state-of-the-art approaches. Under only 5.6% token retention, it preserves 91.46% and 92.06% of the original LLaVA-1.5-7B (Liu et al. 2024a) and LLaVA-NeXT-7B (Liu et al. 2024b) performance, surpassing the best baseline by 0.96% and 1.71%, respectively.
@@
The evolution of large vision-language models (LVLMs) has been driven by the impressive generalization and reasoning abilities of large language models (LLMs) (Touvron et al. 2023; Jiang et al. 2023; Bai et al. 2023; Qwen et al. 2025; Cai et al. 2024), extending their success to the multimodal domain. Recent LVLMs integrate visual and textual modalities by projecting visual inputs into token sequences compatible with language models (Liu et al. 2023; Li et al. 2024; Wang et al. 2024; Bai et al. 2025; Chen et al. 2025b; Zhu et al. 2025), enabling them to leverage the full capacity of LLMs. However, the information density in visual data is often much lower than in textual data, resulting in an excessive number of visual tokens. For instance, LLaVA-1.5 (Liu et al. 2024a) encodes a $\times 336$ 336 image into 576 tokens, while LLaVA-NeXT (Liu et al. 2024b) expands this to 2,880 tokens for high-resolution input. This issue becomes even more critical in video understanding tasks. For example, LongVA (Zhang et al. 2024) transforms 2,000 frames into more than 200K visual tokens, while LongVILA (Chen et al. 2025a) supports inputs exceeding 6,000 frames and produces over a million visual tokens. The inflated visual token sequences place heavy demands on computation and memory, motivating the need for effective token reduction strategies to improve scalability and efficiency across tasks.
@@
Figure 3: Visualization of text-guided visual token retention after shallow-layer pruning (specifically, the second layer) in the LLM. Pale translucent blocks indicate pruned tokens. Retained tokens consistently cluster at the bottom of the image, revealing a positional bias caused by the causal attention in autoregressive LLMs.
@@
To mitigate the computational burden in LVLMs, recent studies have explored reducing the number of visual tokens. Existing approaches to visual token reduction are primarily applied either in the visual encoder or within the LLM. They typically use either text-guided or text-agnostic scoring strategies to determine token importance. Visual encoder pruning approaches perform token reduction before tokens are fed into the LLM. Some approaches rely solely on visual information: FlowCut (Tong et al. 2025) leverages token-level information flow across layers to identify and prune redundant visual tokens in a progressive and structure-aware manner. VisionZip (Yang et al. 2025) and VisPruner (Zhang et al. 2025b) select dominant tokens based on attention scores and compress the rest using similarity-based merging or pruning. VScan (Zhang et al. 2025a) applies global-local scanning. Other works utilize textual guidance at this stage. CDPruner (Zhang et al. 2025c) maximizes instruction-conditioned diversity of retained tokens. SparseVLM (Zhang et al. 2025d) computes token importance via cross-modal attention with text guidance. Although visual encoder pruning benefits from more structured token semantics, it may risk losing fine-grained cues critical for multimodal reasoning. LLM decoding pruning is an alternative direction that performs token reduction within the LLM, where visual and textual tokens are jointly processed. Most of these approaches adopt text-guided token selection. FastV (Chen et al. 2024b) ranks tokens based on their attention from text during generation, and VScan also extends its pruning into LLM layers based on cross-modal attentions. PyramidDrop (Xing et al. 2025) selects visual tokens that are most relevant to the final instruction token, performing progressive pruning across LLM layers. However, these approaches rely on the assumption that visual and textual tokens remain semantically aligned within the LLM, a property that is not guaranteed due to modality entanglement introduced by joint self-attention.
@@
To alleviate the computational overhead in LVLMs, recent studies have explored visual token reduction, either in the visual encoder (Yang et al. 2025; Zhang et al. 2025d) or during the LLM decoding phase (Chen et al. 2024b; Xing et al. 2025). In the latter case, mainstream approaches typically adopt text-guided strategies, using text relevance scores to determine which visual tokens to retain. However, this raises a critical question: Are visual and textual representations truly well-aligned within the LLM layers? Although visual inputs are projected into the textual embedding space before entering the LLM, using text-guided strategies for visual token reduction introduces three forms of cross-modal misalignment, due to the modality-specific biases in how visual and textual information are processed.
@@
In text-guided pruning, the last text instruction token is commonly used to assess the importance of visual tokens. However, due to the causal nature of autoregressive LLMs, this token tends to focus disproportionately on later-positioned visual tokens in the input sequence (Zhang et al. 2025a). Unlike the [CLS] token or average-pooled visual features in bidirectional encoders, the instruction token does not serve as an explicit global aggregator. Figure 3 illustrates this issue, where shallow-layer pruning results show that the retained visual tokens consistently cluster near the end of the input sequence, regardless of their semantic relevance.
@@
Visual token reduction in LVLMs has been explored at two primary stages: either within the visual encoder (Yang et al. 2025; Zhang et al. 2025d, b) or during the LLM decoding phase (Chen et al. 2024b; Xing et al. 2025). We observe that pruning at the encoder stage tends to be more stable, as visual tokens remain semantically coherent and well-structured. However, this may lead to the loss of fine-grained visual details that are crucial for downstream reasoning. In contrast, pruning in the LLM benefits from earlier visual-textual integration, where key visual cues may already be embedded in textual tokens, allowing more informed reduction. Yet as tokens propagate through LLM layers, visual information becomes increasingly entangled and diffused, making it difficult to reliably estimate token importance.
To address the cross-modal misalignment within LLMs and to leverage the complementary strengths of different pruning stages, we propose a training-free, visual-only pruning framework that performs progressive visual token reduction at both the encoder and LLM stages, balancing stability and expressiveness across the model hierarchy, as illustrated in Figure 5.
@@
We partition the LVLM architecture into a sequence of stages $\mathcal{S}={s_{0},s_{1},\dots,s_{N}}$, including both the visual encoder and layers in the LLM decoder. At the end of each stage $s_{n}$, where $n=1,2,\dots,N$, we apply a stage-specific pruning ratio $\lambda_{n}$ to the current set of visual tokens $\mathbf{x}_{V}^{(s_{n})}$, where $|\mathbf{x}_{V}^{(s_{n})}|=\lambda_{n}\cdot|\mathbf{x}_{V}|$. Tokens are ranked based on visual-only attention-based importance scores, and the top-ranked subset is propagated to the next stage $s_{n+1}$. The selection is given by:
@@
To identify the most informative visual tokens, we compute importance scores by reusing the model’s self-attention maps. Specifically, we measure how frequently each token is attended to by visual query tokens, averaged over all query positions. For clarity, we present the single-head attention formulation and omit the pruning stage notation.
@@
FastV (ECCV24) 52.62 57.74 48.43 1540 75.59 68.07 70.51 52.77 46.76 88.45%
PDrop (CVPR25) 57.27 63.51 56.28 1778 82.40 69.56 75.57 56.10 48.95 96.11%
SparseVLM-v1.5 (ICML25) 59.44 65.41 58.69 1789 86.45 68.86 77.01 57.76 50.64 98.64%
VisionZip (CVPR25) 59.25 64.46 57.29 1767 86.39 68.86 76.78 57.26 51.55 98.11%
VisPruner (ICCV25) 59.52 64.91 57.62 1779 86.57 68.52 76.96 57.42 51.46 98.38%
VisionDrop (Ours) 59.99 65.19 58.41 1801 87.23 69.06 77.28 57.81 50.01 98.76%
@@
FastV (ECCV24) 51.83 57.29 49.05 1502 75.55 68.37 66.61 51.09 49.41 87.88%
PDrop (CVPR25) 55.70 62.28 54.32 1656 81.43 69.36 74.21 55.32 49.01 94.04%
SparseVLM-v1.5 (ICML25) 58.43 65.25 58.86 1750 86.37 68.57 76.21 56.61 50.27 97.76%
VisionZip (CVPR25) 57.62 63.40 56.67 1768 84.69 68.82 75.60 56.81 52.02 97.16%
VisPruner (ICCV25) 58.48 63.57 56.67 1743 85.67 69.01 75.74 57.00 52.60 97.53%
VisionDrop (Ours) 58.61 64.52 57.06 1777 85.92 68.52 76.24 57.63 51.06 97.80%
@@
FastV (ECCV24) 48.70 52.30 43.27 1409 67.21 69.36 61.08 49.69 50.55 83.13%
PDrop (CVPR25) 49.51 52.97 43.83 1429 68.42 69.96 63.04 50.80 50.02 84.23%
SparseVLM-v1.5 (ICML25) 53.77 61.27 52.13 1591 80.86 69.61 70.24 53.46 50.35 92.06%
VisionZip (CVPR25) 55.16 62.61 54.20 1687 80.45 69.01 72.42 55.48 52.94 94.62%
VisPruner (ICCV25) 55.59 62.56 54.48 1679 81.29 68.82 72.64 55.83 53.00 94.89%
VisionDrop (Ours) 55.89 62.95 55.10 1698 81.58 69.31 73.16 55.59 52.28 95.22%
Retain 32 Tokens (↓94.4%)
VisionZip (CVPR25) 51.80 58.02 49.78 1592 75.11 68.72 67.12 53.17 52.91 89.92%
VisPruner (ICCV25) 51.98 59.02 50.78 1593 76.44 69.26 67.41 53.47 52.16 90.50%
VisionDrop (Ours) 52.79 60.31 52.91 1572 77.19 69.41 68.55 53.56 52.26 91.46%
Table 1: Performance comparison on LLaVA-1.5-7B (Liu et al. 2024a) under different token retention rates. The best results are marked in bold, and the second-best are underlined.
@@
[…]
Upper Bound, 2880 Tokens (100%)
LLaVA-NeXT-7B 64.27 69.39 62.00 1854 87.54 70.30 81.83 61.40 57.63 100%
Retain 640 Tokens (↓77.8%)
FastV (ECCV24) 56.26 64.97 59.25 1646 83.98 68.27 73.90 54.80 55.00 92.62%
PDrop (CVPR25) 61.80 68.11 61.10 1782 86.44 69.36 79.71 59.44 55.44 97.42%
VisionZip (CVPR25) 61.31 66.14 60.43 1805 87.38 67.82 79.13 60.18 57.32 97.33%
VisionDrop (Ours) 61.70 66.98 60.43 1807 87.85 68.32 79.76 60.45 56.78 97.72%
Retain 320 Tokens (↓88.9%)
FastV (ECCV24) 53.49 61.94 54.99 1495 76.16 66.83 68.24 51.40 52.07 86.82%
PDrop (CVPR25) 57.08 64.46 56.67 1672 79.85 69.91 74.81 54.57 53.66 91.93%
VisionZip (CVPR25) 58.92 63.40 57.40 1712 84.63 67.67 76.32 58.86 56.64 94.26%
VisionDrop (Ours) 59.95 64.91 58.13 1789 85.21 67.97 77.67 59.24 56.97 95.71%
Retain 160 Tokens (↓94.4%)
VisionZip (CVPR25) 55.55 60.09 54.71 1626 79.63 68.12 71.72 56.24 55.95 90.35%
VisionDrop (Ours) 57.02 62.00 55.61 1646 81.88 67.53 73.99 57.52 56.65 92.06%
Table 2: Performance comparison on LLaVA-NeXT-7B (Liu et al. 2024b) under different token retention rates.
@@
We evaluate the proposed VisionDrop on the widely adopted LLaVA-1.5-7B (Liu et al. 2024b) and LLaVA-NeXT-7B (Liu et al. 2024b) for image understanding, as well as Video-LLaVA-7B (Lin et al. 2024) for video understanding. We compare our method with several state-of-the-art visual token reduction baselines, including FastV (Chen et al. 2024b), PyramidDrop (Xing et al. 2025), SparseVLM (Zhang et al. 2025d), VisionZip (Yang et al. 2025), and VisPruner (Zhang et al. 2025b).
@@
We evaluate our method on 9 standard image understanding benchmarks. These include general visual question answering datasets such as GQA (Hudson and Manning 2019), VQAv2 (Goyal et al. 2017), and TextVQA (Singh et al. 2019); robustness and fairness benchmarks including MMBench (Liu et al. 2024c), MMBench-CN (Liu et al. 2024c), MME (Fu et al. 2024) and POPE (Li et al. 2023); as well as specialized datasets like VizWiz (Gurari et al. 2018), which features low-quality images captured by visually impaired users, and ScienceQA (SQA) (Lu et al. 2022), which focuses on scientific reasoning. Additionally, we evaluate our method on 3 video understanding benchmarks, TGIF (Jang et al. 2017), MSVD (Chen and Dolan 2011), and MSRVTT (Xu et al. 2016).
@@
We adopt the standard inference settings provided in the official implementations of each LVLM. For progressive visual token pruning, we divide the model into five stages: the first stage concludes at the output of the visual encoder, while the remaining four end at decoding layers $l=8$, 161616, 242424, and the final decoding layer of the LLM. The number of visual tokens is gradually reduced across these stages. In the first stage, all original visual tokens are retained. For each reduction configuration, the retention number in the second stage is set to 1.5×1.5$\times$1.5 × the final count for image understanding and 3×3$\times$3 × for video understanding, and tokens are pruned proportionally in subsequent stages to meet the overall budget.
Results and Analyses
Results on LLaVA-1.5.